There is this question on my homework and it says: write and equation in standard form of the line that passes through (7,2) and has a y- intercept of -4
I know that standard form is Ax=By=C and i'm thinking that you would make it (7,2) (0,4) and then it would go into Y2-Y1/X2-X1 form. And end up with -6/-7 and now I and don't know what to do from here.

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

There is this question on my homework and it says: write and equation in standard form of the line that passes through (7,2) and has a y- intercept of -4
I know that standard form is Ax=By=C [it's Ax + By = C]and i'm thinking that you would make it (7,2) (0,4) and then it would go into Y2-Y1/X2-X1 form. And end up with -6/-7 and now I and don't know what to do from here.
-------------
The points are (7,2) and (0,-4)
slope m = diffy/diffx = -6/-7 = 6/7
----
You have the slope and the y-intercept, --> eqn is
y = (6/7)x + 4 (slope-intercept form)
Multiply by 7
7y = 6x + 28
6x - 7y = -28 (standard form)

Answer by richwmiller(17219)

y = (6/7)x - 4
7y=6x-28
6x-7y=28